Problem (8 pts) a) What mass of potassium nitrate (M 101.1 gmol) is required to prepare 300 mL of 0.1 M solution? bo What will be the concentration of potassium ions if extra 10 g of potassium sulfate 174.3 g/mol) are added? Assume (M that the volume of the salt is negligible.
Answer #1

(a)

Molarity = (mass / molar mass) * (1000 / Volume of solution in L)

0.1 = (mass / 101.1) * (1000 / 300)

Mass of KNO3 = 3.03 g.

KNO3 (aq.) --------------> K+ (aq.) + NO3- (aq.)

[K+] = [KNO3] = 01. M        initial

Intial moles of K+ ions = 0.1 * 300 / 1000 = 0.0300 mol

(b)

Moles of K2SO4 added = 10 / 174.3 = 0.0574 mol

K2SO4 (s) --------------> 2 K+ (aq.) + SO42- (aq.)

Moles K+ = 2 * moles of K2SO4 = 2 * 0.0574 = 0.115 mol

Therefore, total moles of K+ ions = 0.0300 + 0.115= 0.145 mol

Therefore, new [K+] = 0.145 * 1000 / 300 = 0.483 M